sofiaarbo:

3. Adipic acid, a raw material for nylon, is made industrially by the oxidation of cyclohexane. 5 O2 + 2 C6H12 ----- > 2 C6H10O4 + 2 H2O cyclohexane special conditions adipic acid (a) How many moles of oxygen would be needed to make 40.0 moles of adipic acid by this reaction? (b) If 164 grams of cyclohexane is used, what is the theoretical yield of adipic acid in moles? In grams?

Vocaloid:

the equation is already balanced so let's jump into the stoich a) we want to convert 40.0 moles of adipic acid to moles oxygen. we can use the coefficients of the equation \[(\frac{ 40.0~moles~Adipic~acid }{ })(\frac{ 5~moles~oxygen }{ 2~moles~Adipic~Acid })\] to convert to moles oxygen

Vocaloid:

b) similar logic with converting 164g of cyclohexane to adipic acid cyclohexane molar mass is 84.16g/mol \[\frac{ 164g~cyclohexane }{ }(\frac{ 1~mol~cyclohexane }{ 84.16g~cyclohexane })(\frac{ 2~mol~adipic~acid }{ 2~mol~cyclohexane })\] to get to moles adipic acid, stop here. but since it also asks for grams of adipic acid, multiply by the molar mass of adipic acid (146.14g/mol) please let me know if you have any additional questions about this work, or just want to check your answer

sofiaarbo:

no, that is it. Thank you very much :)